Find the two complementary angles whose one angle measures 5 times greater than the other less 20 degrees.
Answer by stanbon(75887)

Find the two complementary angles whose one angle measures 5 times greater than the other less 20 degrees.
-----
Let the complementary angles be "x" and "90-x".
----
Equation:
(90-x) = 5x - 20
70 = 6x
x = 70/6 = 11 2/3 degrees
---
90-x = 78 1/3 degrees
========================
Cheers,
Stan H.
========================